# Koopman Operators in Robot Learning

### v1

@@ IV-A2 Continuous-Time Case
    where $\cdot$ and $\nabla$ represent the dot product and the gradient operator respectively. $\mathcal{L}_{G}$ is often referred to as the Koopman generator.666 The definition of Koopman generator can be slightly relaxed to hold on a dense subset of $\mathcal{F}$ [33, Chapter 1]. By convention, we define the eigenfunctions of the Koopman semi-group slightly differently from the typical definition. A nonzero function $\phi\in\mathcal{F}$ is an eigenfunction of the semi-group $\{\mathcal{K}^{t}\}_{t\in\mathbb{R}_{\geq 0}}$ with eigenvalue $\lambda_{\phi}\in\mathbb{C}$ if
    $\mathcal{K}^{t}\phi=e^{\lambda_{\phi}t}\phi,\quad\forall t\in\mathbb{R}_{\geq 0}\;.$ (14)
    For strongly continuous semi-group of operators, given the Koopman generator, (14) leads to the equivalent definition
    $\mathcal{L}_{G}\phi=\lambda_{\phi}\phi\;.$ (15)
    It is important to observe the similarity between (15) and (8). The concept of invariant subspaces for the continuous case is then analogous to the discrete-time case.
    IV-B Extended Dynamic Mode Decomposition (EDMD)
    In practice, we have access to discrete measurement data of the system to obtain a finite-dimensional approximated matrix of the infinite-dimensional Koopman operator. Extended Dynamic Mode Decomposition (EDMD) approach [25] is among the most widely-used approximation methods.
    Consider data matrices taken from system (6) as
    $X=[x_{1},x_{2},\ldots,x_{M}]\;,$
    $Y=[y_{1},y_{2},\ldots,y_{M}]\;,$
    where $y_{i}=T(x_{i})$ for all $i\in\{1,\dots,M\}$; we can also use the state history $Y=[x_{2},\dots,x_{M},x_{M+1}]$. We need a dictionary of functions that lift state variables to the higher-dimensional space where the observable dynamics is approximately linear. Define the dictionary $\mathbf{\Psi}(x_{m})=[\psi_{1}(x_{m}),\ldots,\psi_{N}(x_{m})]^{T}$. Then, the Koopman operator can be approximated by minimizing the total residual between snapshots,777 For ease of exposition our notation is transpose of the notation in [25].
    $J=\frac{1}{2}\sum_{m=1}^{M}\|(\mathbf{\Psi}\left(x_{m+1}\right)-K\mathbf{\Psi}\left(x_{m}\right)\|^{2}\enspace.$ (16)
    This can be solved as a least-square problem, yielding
    $K_{\operatorname{EDMD}}=\bm{A}\bm{G}^{\dagger},\enspace\text{where }\begin{cases}\bm{G}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{\Psi}(x_{m})\mathbf{\Psi}^{*}(x_{m}),\\ \bm{A}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{\Psi}(x_{m+1})\mathbf{\Psi}^{*}(x_{m}),\end{cases}$ (17)
    with being the pseudoinverse and the conjugate transpose.
[…]
    with the closed-form solution
    The approximation error of the predictors in (20)-(21) depends on the selected dictionary of functions $\mathbf{\Psi}$. More precisely, the error depends on how close $\operatorname{span}(\mathbf{\Psi})$ is to be invariant under the Koopman operator. To understand this, it is important to observe that $K_{\operatorname{EDMD}}$ does not capture the Koopman operator itself but it encodes the projection of the operator’s action on $\operatorname{span}(\mathbf{\Psi})$, i.e. EDMD approximates the following linear operator
    $\mu_{X}=\frac{1}{M}\sum_{i=1}^{M}\delta_{x_{i}}\;,$ (23)
    with $\delta_{x_{i}}$ the Dirac measure at point $x_{i}$.
    In general, to observe the asymptotic convergence phenomena of EDMD to the Koopman operator described in [35], the dimension of the dictionary and the number of data points must be sufficiently large. Moreover, without a system’s model, it is not possible to estimate a lower bound on the dictionary’s dimension to achieve a predetermined level of accuracy. In addition, if the dictionary sequence is chosen based on a given basis for the space $\mathcal{F}$ (that is not chosen based on the system’s knowledge), the dimension of the dictionary required to achieve a high accuracy level might be extremely large. Therefore, for practical applications, it is imperative to design or learn dictionaries based on information available from the system and/or data to achieve a reasonable accuracy on relatively low-dimensional subspaces. We refer the reader to Section V-B for more information on such methods. One last observation is that these convergence results only hold for systems without control inputs. Further discussion on these can be found in Section VII.
    Remark IV.1
    IV-C Koopman Approximation for System with Inputs
    The Koopman operator theory and EDMD were originally proposed without the consideration of inputs. To generalize the Koopman theory to allow for control inputs, the corresponding dynamical systems can be extended from (6) and (10) as
    $x_{t+1}=T_{u}(x_{t},u_{t})\text{ and }\dot{x}=G_{u}(x,u)\;,$ (24)
    with $u\in\mathbb{R}^{N_{u}}$. For convenience, we only discuss the case of discrete-time systems in the following (the treatment for continuous-time systems is similar). Define a set of observables $g(x,u)$ that are functions of both the state and the input. The Koopman operator acting on the space of those observables is expressed as
    $\mathcal{K}g(x_{t},u_{t})=g\circ T_{u}=g(T_{u}(x_{t},u_{t}),u_{t+1})\;.$ (25)
    Because (24) only prescribes how the state evolves, this formulation is, in general, ill-posed as it treats the state and the input as being the same when they are not. However, this formulation can work in specific cases [41]:
    1. 1.
      If the input is a time-invariant constant forcing signal, i.e. $u_{t+1}=u_{t}=c$, (25) can be rewritten as $\mathcal{K}g(x_{t},c)=g(T_{u}(x_{t},c),c)$.
    2. 2.
      If the input is generated as a function of states, i.e. via closed-loop control so that $u_{t}=f_{x}(x_{t})$, (25) can be simplified to contain only the state as $\mathcal{K}\hat{g}(x_{t})=\hat{g}(\hat{T_{u}}(x_{t}))$, where $\hat{T_{u}}$ is a function of $T_{u}$ and $f_{x}$, and $\hat{g}$ is a function of $g$ and $f_{x}$. Thus, a known closed-loop control law can reduce the complexity of the Koopman expression.
    3. 3.
      If the input follows its dynamics, i.e. $u_{t+1}=f_{u}(u_{t})$, (25) can also be simplified as $\mathcal{K}g(\hat{x}_{t})=g(\hat{T_{u}}(\hat{x}_{t}))$ by treating the inputs as states, i.e. $\hat{x}=[x;u]$.
    When the input is generated from exogenous forcing, which is common in robotics, then (25) keeps its general form. Approaches to consider in this case include approximating the action of the inputs on the evolution of observables linearly (i.e. $g\circ T_{u}=\mathcal{K}g(x_{t})+Bu_{t}$) or bilinearly (i.e. $g\circ T_{u}=\mathcal{K}g(x_{t})+Bu_{t}+g(x_{t})^{T}Fu_{t}$). Through the Koopman canonical transform [42, 29], control-affine systems $x_{t+1}=T(x_{t})+\sum_{i}h_{i}(x_{t})u_{t}^{i}$ can be transformed into a bilinear form.
    Dealing with control system (24) is significantly more difficult than system (6). The difficulty is rooted in the difference between the input and the state. The state is a fundamental property of the system and evolves according to its dynamics, while the input in an open loop system is not known a-priori, and its choice can significantly alter the system’s behavior. Below, we discuss two different ways to deal with this fundamental issue.
    Considering All Infinite Input Sequences
@@ IV-C Koopman Approximation for System with Inputs
    The work in [43] tackles the aforementioned issue by considering the system’s behavior over all possible input sequences. Formally, consider the space $l(\mathcal{U})$ comprised of all possible infinite input sequences $\{u_{i}\}_{i=0}^{\infty}$ with $u_{i}\in\mathcal{U}$ and let $S:l(\mathcal{U})\to l(\mathcal{U})$ denote the left-shift operator on $l(\mathcal{U})$, mapping the sequence $\{u_{i}\}_{i=0}^{\infty}$ to $\{u_{i}\}_{i=1}^{\infty}$. Now, consider the extended state $x_{s}:=(x,u_{s})\in\mathcal{X}\times l(\mathcal{U})$ and define the dynamical system
@@
    $z^{+}=Az+Bu\;,$ (28)
    where $z$ is the lifted state starting from the initial condition $z_{0}=\mathbf{\Psi}(x_{0})$ (here $\mathbf{\Psi}:\mathcal{X}\to\mathbb{R}^{N_{\mathbf{\Psi}}}$ is the lifting map). One can estimate the matrices $A$ and $B$ in (28) with an EDMD-like method. However, it is crucial to keep in mind that since the model does not consider infinite-input sequences, it does not generally capture all the information of the operator $\mathcal{K}_{L}$. Even if the dimension of the lifted state goes to infinity, one cannot generally conclude convergence of the lifted linear model trajectories to the trajectories of the nonlinear system (24); see also [43, Discussion after Corollary 1].
    Koopman Control Family
    An alternative Koopman-based approach to model the system (24) does not rely on infinite input sequences and it is easier to use for finite-dimensional approximations [44]. The central idea is that one can represent the behavior of (24) by a family of systems in the form of (6) generated by setting the input in (24) to be a constant signal. Consider the family of constant input dynamics $\{T_{\hat{u}}\}_{\hat{u}\in\mathcal{U}}$ defined by
@@
    Similarly to the case without control inputs where Koopman-invariant subspaces lead to a finite-dimensional linear form (cf. (9)), the models on common invariant subspaces of the Koopman control family are all in the following form (termed “input-state separable model”, cf. [44, Theorem 4.3])
@@
    where $\bm{\Phi}:\mathcal{X}\to\mathbb{C}^{N_{\bm{\Phi}}}$ is the lifting function and $\mathcal{A}:\mathcal{U}\to\mathbb{C}^{N_{\bm{\Phi}}\times N_{\bm{\Phi}}}$ is a matrix-valued function. Note that the input-state separable model is linear in the lifted state but nonlinear in the input. This is because, in an open-loop system, the input does not abide by predefined dynamics. Therefore, one cannot use the structure of the Koopman operator to represent the nonlinearity in the input as a linear operator. Interestingly, the popular lifted linear, bilinear, and linear switched models (e.g., [45]) are all special cases of the input-state separable form (cf. [44, Lemmas 4.4-4.5]). In case there is no suitable common invariant subspace, one can approximate the action of the Koopman Control Family on any given subspace via orthogonal projections. See [44] for theoretical analysis, data-driven learning methods, and accuracy bounds.
    V Implementation Details
    Having discussed the overall implementation as well as the fundamental theoretical results underlying the Koopman operator in robotics, we now take a deeper look into important implementation steps covering data collection, model estimation (with a specific focus on the selection of lifting functions), and controller design. Here, we present a comprehensive array of specific approaches and engage in detailed discussions regarding various options within each step.
    V-A Data Collection
    As in all data-driven algorithms, the type and amount of (training) data to collect play an important role in model accuracy and control efficacy. Optimizing the data collection process in Koopman-based methods is an active research direction [46, 47]. Methods focus on the collection of both offline and online measurements.
    One of the most common methods to collect data is random selection [48], i.e. generating measurements by propagating the system with random initial conditions and input signals. To enrich the diversity of the data, one should select initial conditions and inputs over the entire operating range of the robot [46, 27]. An enclosed area might be required for safety [49]. The random collection approaches are more widely used for soft robots [50, 51, 52] as soft parts do less harm to their surroundings if an aggressive command is selected by chance.
    However, generating data with random inputs can compromise safety, especially for rigid robots. One attempt to solve this challenge is to design a baseline controller, which can be open-loop and naive but is improved over time. For example, [31] utilizes the controller obtained from the previous episode to generate training data used for the current episode. By doing that, we can obtain observations in a safer way and with more valuable information. Another way that emphasizes the effectiveness of data is to account for the value of information, i.e. add constraints on information richness when solving the optimization problem. These ideas are closely connected to active learning [30], which we will discuss in Section V-C2.
    For methods that rely on extracting dynamics from data, it is important to consider the stochastic case, when the measurements used for estimation are noisy, as the efficiency and quality of data have a significant impact on the performance of the estimated model and hence the controller design. To quantify and emphasize the effect of noise, [53] presents a procedure for deriving loose and tight bounds of the prediction error using Dynamic Mode Decomposition (DMD). The general version for EDMD is explored in [54]. Once the explicit prediction error due to noisy measurements is derived, this work describes how the proposed algorithm and the resulting prediction error can be utilized in controller design, providing robustness guarantees for the perturbed systems’ performance. Another approach [48] to minimize the effect of modeling uncertainties and external disturbance is to include the usage of a Kalman filter. First, an augmented model is derived to include the disturbance model in the ordinary Koopman structure, and then a Kalman filter is adapted to estimate the observables and disturbances simultaneously. We delve into a comprehensive discussion about understanding and dealing with the impact of measurement noise when estimating the Koopman operator through EDMD and DMD.
    V-A1 EDMD and DMD with Measurement Noise
    Here we discuss methods that deal with the effect of measurement noise in data following [55]. To better understand the effect of measurement noise in EDMD, we reformulate (18) as
    $\underset{K,\Delta}{\text{minimize}}$ $\|\Delta\|_{F}^{2}$
    subject to $\mathbf{\Psi}(Y)^{T}+\Delta=\mathbf{\Psi}(X)^{T}K^{T}\;.$ (31)
    This reformulation 888 The notation in [55] is the transpose of the notation employed here. immediately reveals two issues when data matrices are noisy.
    • Lack of correction in $\mathbf{\Psi}(X)$: Based on (V-A1), EDMD finds a correction for $\mathbf{\Psi}(Y)$ with minimum Frobenius norm subject to the optimization constraints. However, EDMD does not provide any correction for $\mathbf{\Psi}(X)$. If $\mathbf{\Psi}(X)$ is noisy, the EDMD solution will be incorrect. This is a limitation inherent to least-squares methods where noise in both dictionary matrices $\mathbf{\Psi}(Y)$ and $\mathbf{\Psi}(X)$ can lead to inconsistent solutions [56].
    • Noise distortion after applying the nonlinear map $\mathbf{\Psi}$: Since we measure the data $X,Y$ and pass them through $\mathbf{\Psi}$, noise in $X,Y$ gets distorted. (Even if the noise in $X,Y$ is zero-mean and i.i.d, distorted noise in $\mathbf{\Psi}(Y),\mathbf{\Psi}(X)$ is in general neither.) Corrections to $\mathbf{\Psi}(Y),\mathbf{\Psi}(X)$ are needed to counter the noise distortion.
    To address these issues, one should add a correction term for $\mathbf{\Psi}(X)$ and introduce additional terms canceling the noise’s distortion. This leads to
    $\underset{K,\Delta_{1},\Delta_{2}}{\text{minimize}}$ $\sum_{i=1}^{M}\big{\|}[\Delta_{1},\Delta_{2}]_{i}\;C_{i}^{-1/2}\big{\|}_{2}^{2}$
    subject to $\;\mathbf{\Psi}(Y)^{T}-B_{\mathbf{\Psi}(Y)^{T}}+\Delta_{2}$
    $=(\mathbf{\Psi}(X)^{T}-B_{\mathbf{\Psi}(X)^{T}}+\Delta_{1})K^{T}\;,$ (32)
    where $[\Delta_{1},\Delta_{2}]_{i}$ denotes the $i$th row of the matrix created by concatenating $\Delta_{1},\Delta_{2}$ side by side. Here, $B_{\mathbf{\Psi}(X)^{T}},B_{\mathbf{\Psi}(Y)^{T}}$ account for the bias created by noise distortion in $\mathbf{\Psi}(X)^{T},\mathbf{\Psi}(Y)^{T}$, and $C_{i}^{-1/2}$’s account for the covariance distortion (for simplicity, we do not provide the precise expressions, but explicit formulas of the bias and covariance corrections can be found in [55]). The problem (V-A1) is an element-wise weighted total least squares optimization [57, 58]. Under appropriate conditions on data-sampling and dictionary functions, (V-A1) applied on noisy data provides a weakly consistent estimator of the EDMD’s solution applied on noise-free data (see e.g., [57, 55]).
    The measurement noise problem in DMD can be addressed more directly since there is no lifting via a nonlinear dictionary, so noise distortion does not occur. Hence, instead of solving (V-A1) one only needs to form a total least squares problem [59, 60] as
    $\underset{K,\Delta_{1},\Delta_{2}}{\text{minimize}}$ $\big{\|}[\Delta_{1},\Delta_{2}]\big{\|}_{F}^{2}$
    subject to $\;Y^{T}+\Delta_{2}=(X^{T}+\Delta_{1})K^{T}\;,$ (33)
    Under some conditions, (V-A1) provides a strongly consistent estimator for the DMD’s solution on noise-free data [61].
    An alternative approach to deal with noise in DMD is to consider both forward and backward dynamics. Given that for a nonsingular linear system, the forward and backward dynamics are inverse of each other, the work in [59] proposes the use of forward and backward DMD on the data to account for the effect of noise that is assumed to be additive to data, normally distributed, and with zero mean. In this case, one can approximate the correct DMD matrix as
    $\tilde{K}_{\text{DMD}}\approx\big{(}A_{\text{DMD}}^{f}A_{\text{DMD}}^{b}\big{)}^{-\frac{1}{2}}\;,$
    where $A_{\text{DMD}}^{f}$ and $A_{\text{DMD}}^{b}$ are the debiased solutions of DMD on the noisy data forward and backward in time. Alternatively, [62] proposes a different method to consider the forward and backward dynamics by solving
[…]
    This optimization problem aims to find the matrix $K$ to account for forward and backward dynamics. The problem is highly non-convex and hence difficult to solve: the work in [62] provides several algorithms to solve it efficiently.
    V-B Model Estimation — Lifting Functions Design
    The lifting functions are crucial because they serve as the basis for constructing a linear approximation of the (nonlinear) system’s state evolution. A poor choice of the lifting functions can significantly impact the estimation accuracy of the Koopman operator and the higher-dimensional linear dynamics.
    Existing examples regarding the design of the lifting functions fall under several main directions. The first one is empirical selection. For example, discontinuous spectral elements can make the observation matrix block diagonal [25], Hermite polynomials are best suited to problems where data are normally distributed, and radial basis functions are effective for systems with complex geometry. However, empirical approaches can be time- and effort-demanding and cannot guarantee generalization to and efficiency in new cases. Another direction is mechanics-inspired selections since robotic systems have certain characteristic properties, e.g., degrees of freedom, configuration spaces, and workspaces, that can be acquired without knowledge of their exact dynamical models. These properties reveal fundamental information about system states and dynamics and can provide intuition on how to select lifting functions required for Koopman operator approximation [63]. A third approach is to synthesize basis functions out of higher-order derivatives of the states (in the continuous time) [64] or past measurements (commonly referred to as Hankel Koopman models) [65], which can be used for completely unknown dynamics. Lastly, one can compare and optimize over multiple sets of lifting functions to find a proper basis set, but the question of how to select these sets efficiently remains open.
    Next, we describe in detail data-driven methods for the identification of dictionaries/subspaces that lead to accurate Koopman approximations. These methods have in common their reliance on the notion of Koopman-invariant subspace and the fact that the restriction of the Koopman operator to such subspaces is exact. They are highly connected to the convergence analysis introduced earlier in Section IV-B.
    V-B1 Optimization-based Methods
    Learning in original coordinates: Available data can be used to learn an appropriate dictionary. The work [66] takes advantage of the eigenfunction evolution on non-recurrent sets and provides an optimization-based approach to directly learn Koopman eigenfunctions which naturally span invariant subspaces, leading to accurate long-term prediction. The work [67] uses a sparsity-promoting approach to derive the learned subspace to being close to invariant under the Koopman operator. Another popular approach is the use of neural networks or other parametric families (such as polynomials, Fourier, or radial-basis functions) to learn appropriate dictionaries (while fixing the state variables as a part of the dictionary) through the optimization
    $\underset{K\in\mathbb{R}^{M\times M},\mathbf{\Psi}\in\text{PF}}{\text{minimize}}\|\mathbf{\Psi}(Y)-K\mathbf{\Psi}(X)\|_{F}\;,$ (34)
    where PF is the parametric family of choice. This optimization can be solved simultaneously in both variables $K$ and $\mathbf{\Psi}$ [68], in a sequential manner in each variable [69], or one can simply use the closed-form solution (19) of EDMD [70], to write
    $\underset{\mathbf{\Psi}\in\text{PF}}{\text{minimize}}\|\mathbf{\Psi}(Y)-\mathbf{\Psi}(Y)\mathbf{\Psi}(X)^{\dagger}\ \mathbf{\Psi}(X)\|_{F}\;.$ (35)
    Note that the cost function in (35) is the residual error of EDMD at the optimal solution. It is important to note that (34)-(35) are prone to over-fitting, therefore regularization may be necessary. More importantly, optimization problems (34)-(35) are solved over a parametric family of functions (not just the EDMD matrix); therefore, they are generally non-convex and finding globally optimal solutions is not guaranteed.

### v2

@@ V-A Koopman Operator Theory for Continuous-Time Systems
where $\cdot$ and $\nabla$ represent the dot product and the gradient operator respectively. $\mathcal{L}_{G}$ is often referred to as the Koopman generator.666 The definition of Koopman generator can be slightly relaxed to hold on a dense subset of $\mathcal{F}$ [160, Chapter 1].
V-B Koopman Operators for Systems with Control Inputs
Consider the control system
$x_{t+1}=T_{u}(x_{t},u_{t})\;,$ (21)
with state space $\mathcal{X}\subseteq\mathbb{R}^{N_{x}}$ and input space $\mathcal{U}\subseteq\mathbb{R}^{N_{u}}$. Dealing with systems in (21) from a Koopman operator perspective is significantly more difficult compared to systems without input. This challenge arises from the fundamental distinction between states and inputs: while the state is an intrinsic property of the system that evolves according to its internal dynamics, the input—especially in open-loop systems—is not known a priori and can significantly influence the system’s evolution. In the following, we present two theoretical perspectives that aim to address this foundational issue and offer deeper insights into how inputs can be integrated into Koopman-based models.
V-B1 Considering All Infinite Input Sequences
@@
The work in [106] tackles the aforementioned issue by considering the system’s behavior over all possible input sequences. Formally, consider the space $l(\mathcal{U})$ comprised of all possible infinite input sequences $\{u_{s}(i)\}_{i=0}^{\infty}$ with $u_{s}(i)\in\mathcal{U}$ and let $S:l(\mathcal{U})\to l(\mathcal{U})$ denote the left-shift operator on $l(\mathcal{U})$, mapping the sequence $\{u_{s}(i)\}_{i=0}^{\infty}$ to $\{u_{s}(i)\}_{i=1}^{\infty}$. Now, consider the extended state $x_{s}:=(x,u_{s})\in\mathcal{X}\times l(\mathcal{U})$ and define the dynamical system
@@
$z^{+}\approx Az+Bu\;,$ (24)
where $z$ is the lifted state starting from the initial condition $z_{0}=\mathbf{\Psi}(x_{0})$ (here $\mathbf{\Psi}:\mathcal{X}\to\mathbb{R}^{N_{\mathbf{\Psi}}}$ is the lifting map). One can estimate the matrices $A$ and $B$ in (24) with an EDMD-like method. However, it is crucial to keep in mind that since the model does not consider infinite-input sequences, it does not generally capture all the information of the operator $\mathcal{K}_{L}$. Even if the dimension of the lifted state goes to infinity, one cannot generally conclude convergence of the lifted linear model trajectories to the trajectories of the nonlinear system; see also [106, Discussion after Corollary 1].
V-B2 Koopman Control Family
An alternative Koopman-based approach to model the system does not rely on infinite input sequences and it is easier to use for finite-dimensional approximations [36]. The central idea is that one can represent the behavior of (21) by a family of systems in the form of (1) generated by setting the input in (21) to be a constant signal. Consider the family of constant input dynamics $\{T_{\hat{u}}\}_{\hat{u}\in\mathcal{U}}$ defined by
@@ V-B2 Koopman Control Family
Similarly to the case without control inputs where Koopman-invariant subspaces lead to a finite-dimensional linear form, the models on common invariant subspaces of the Koopman control family are all in the following form (termed “input-state separable model”, cf. [36, Theorem 4.3])
@@
where $\bm{\Phi}:\mathcal{X}\to\mathbb{C}^{N_{\bm{\Phi}}}$ is the lifting function and $\mathcal{A}:\mathcal{U}\to\mathbb{C}^{N_{\bm{\Phi}}\times N_{\bm{\Phi}}}$ is a matrix-valued function. Note that the input-state separable model is linear in the lifted state but nonlinear in the input. This is because, in an open-loop system, the input does not abide by predefined dynamics. Therefore, one cannot use the structure of the Koopman operator to represent the nonlinearity in the input as a linear operator. Interestingly, the popular lifted linear, bilinear, and linear switched models (e.g., [161]) are all special cases of the input-state separable form (cf. [36, Lemmas 4.4-4.5]). In case there is no suitable common invariant subspace, one can approximate the action of the Koopman Control Family on any given subspace via orthogonal projections. See [36] for theoretical analysis, data-driven learning methods, and accuracy bounds.
V-C Lifting Function Construction
Here, we describe methods for the identification of dictionaries that lead to accurate Koopman approximations. The accuracy of such approximations is directly related to the quality of the underlying subspace (the dictionary’s span), specifically how close it is to being invariant under the Koopman operator.
V-C1 Optimization-based Methods
As introduced in the previous section, available data can be leveraged to learn a suitable dictionary of observables using neural networks or other parametric function families. This is typically achieved by minimizing the EDMD’s residual error, formulated by combining equations (6) and (7) as follows:
$\underset{\mathbf{\Psi}}{\text{minimize}}\|\mathbf{\Psi}(Y)-\mathbf{\Psi}(Y)\mathbf{\Psi}(X)^{\dagger}\ \mathbf{\Psi}(X)\|_{F};/.$ (27)